On October 28, 2015, the DJIA opened at 12,608.56. The divisor at that time was 0.132340204. In October 2015, Goldman Sachs was the highest priced stock in the DJIA and Cisco was the lowest. Assume the closing price for Goldman Sachs on October 27, 2015, was $246.48, and the closing price for Cisco was $12.78. Suppose the next day the other 29 stock prices remained unchanged and Goldman Sachs increased 4.4 percent. What would the new DJIA level be? Now assume only Cisco increased by 4.4 percent. Find the new DJIA level.
Answer #1

New DJIA=Opening Level+Increase in price/divisor
What would the new DJIA level be?
=12608.56+246.48*4.4%/0.132340204=12690.50879

Now assume only Cisco increased by 4.4 percent. Find the new DJIA level.
=12608.56+12.78*4.4%/0.132340204=12612.80905